Circular No. 8089

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SATELLITES OF JUPITER Further to IAUC 8088, S. S. Sheppard reports the discovery of four

additional satellites of Jupiter, found on CCD images taken at Mauna Kea during Feb. 6–Mar. 7. The astrometry, orbital elements by B. G. Marsden, and ephemerides appear on MPEC 2003-E29: S/2003 J 9, e = 0.27, i = 164^{o}, P = 684 days, H = 17.2; S/2003 J 10, e = 0.21, i = 164^{o}, P = 768 days, H = 16.7; S/2003 J 11, e = 0.22, i = 164^{o}, P = 682 days, H = 16.8; S/2003 J 12, e = 0.38, i = 146^{o}, P = 533 days, H = 17.2. In revision to IAUC 8087, further calculations (also given on MPEC 2003-E29) suggest alternative retrograde orbits for S/2003 J 1 (e = 0.34, i = 163^{o}, P = 783 days) and for S/2003 J 6 (e = 0.16, i = 156^{o}, P = 618 days).

SUPERNOVAE 2003bo, 2003bp, AND 2003bq Three apparent supernovae have been found on unfiltered CCD images

— SN 2003bo by W. M. Wood-Vasey et al. (cf. IAUC 8088; Palomar NEAT images), SN 2003bp by D. Hutchings and W. Li (LOTOSS, via KAIT images; cf. IAUC 8086), and SN 2003bq by T. Boles (cf. IAUC 8086).

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003bo Feb. 19.41 9^{h}47^{m}43.^{s}41 +36^{o}18′21′′.8 19.6 8′′.4 W
2003bp Mar. 6.2 8 27 27.81 +17 17 14.5 17.8 17′′.6 E, 11′′.6 N
2003bq Mar. 6.95 6 42 50.12 +41 25 19.9 18.7 20′′.1 W, 4′′.3 N

Additional approximate magnitudes from the respective discoverers (those denoted ‘NEAT’ are from Wood-Vasey): SN 2002bo, 2002 Nov. 24.52, [20.2; 2003 Mar. 3.28, 19.5. SN 2002bp in NGC 2596, Feb. 19.2, [19.0; 28.3, 18.2 (NEAT); Mar. 6.1, 17.8 (NEAT); 7.2, 17.8. SN 2003bq in UGC 3513, Feb. 17, [19.5; Mar. 7.791, 18.7. SN 2003bq is also absent from Palomar Sky Survey red and blue plates. SN 1998fa also appeared in UGC 3513.

COMET C/2002 V1 (NEAT) Total visual magnitude estimates: Feb. 24.43 UT, 2: (M. Mattiazzo,

Wallaroo, South Australia, 25×100 binoculars; bright twilight); 25.10, 2.9: (J. Shanklin, Rothera, Antarctica, 10×50 binoculars); 28.12, 4.4 (Shanklin); Mar. 1.90, 4.0 (W. Souza, Mairipora, Brazil, 11×80 binoculars); 4.94, 5.5 (G. R. Hickel, São José dos Campos, Brazil, 16×50 binoculars).

2003 March 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green